OpenStudy (anonymous):

Solve the following system of equations. Please show me how to do it in Substitution method. x - y = 10 2x + y = 2

jimthompson5910 (jim_thompson5910):

solve the first equation for x x - y = 10 x - y+y = 10+y x = 10+y

jimthompson5910 (jim_thompson5910):

Then replace x in the second equation with 10+y (since x = 10+y) So 2x + y = 2 turns into 2(10+y) + y = 2

jimthompson5910 (jim_thompson5910):

now solve 2(10+y) + y = 2 for y once you have the value for y, you can plug it into x = 10+y to get x

OpenStudy (anonymous):

Im gettin it so far.

jimthompson5910 (jim_thompson5910):

ok great, tell me what you get when you solve for y

OpenStudy (anonymous):

We have x−y=10 and 2x+y=2 Let's add both the equations, this will eliminate y x−y+2x+y=12 x−y+2x+y=12 3x=12 divide both sides by 3 we get x=4 Now use first equation x−y=10 , let's substitute x=4 here 4−y=10 subtract both sides by 4 , we get −y=6 we get y=−6 Here we have used elimination method because it was easy to apply here. We just added the two equations to eliminate y and were able to find x

OpenStudy (anonymous):

Ohh okay, the problem says to use the easier way of solving so I though Substitution was easier.

jimthompson5910 (jim_thompson5910):

it depends on the problem and which one you think is easier. Some find elimination easier, while others find graphing easier.

OpenStudy (anonymous):

Oh okay so y = -6 ?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

Okay I got it, thanks for the help. :)

jimthompson5910 (jim_thompson5910):

good so if y = -6, then x = ???

OpenStudy (anonymous):

np(: and @jim_thompson5910 was right too(:

jimthompson5910 (jim_thompson5910):

this question is directed toward 24HourDiet no offense seiga (It looks like you know this stuff well, which is awesome)

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

Both of yall are awesome lmao.

jimthompson5910 (jim_thompson5910):

x =10+y x = 10 + (-6) x = 10 - 6 x = 4 So you are correct

OpenStudy (anonymous):

lol not that awesome...